Given the following information, determine the beta coefficient for Stock L that is consistent with equilibrium: = 10.5%; rRF = 2.5%; rM = 8.5%. Round your answer to two decimal places.

ANSWER = _____   

PLEASE SHOW WORK PLEASE. SHOW HOW IN CALCULATOR OR BASIC MATH STRUCTURE IF POSSIBLE. THANK YOU
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

10.5=2.5+Beta*(8.5-2.5)

(10.5-2.5)=Beta*(8.5-2.5)

Beta=(10.5-2.5)/(8.5-2.5)

=1.33(Approx).